# The $\Lambda_{b}\to\Lambda$ transition form factors in perturbative QCD approach

### v1

@@ Abstract
In this work, we investigate the $\Lambda_{b}\to\Lambda$ transition form factors in the perturbative QCD (PQCD) approach, incorporating higher-twist light-cone distribution amplitudes (LCDAs). The resulted form factors show that higher-twist LCDAs are dominant numerically. By combining our PQCD predictions at low-$q^{2}$ with lattice QCD results at high-$q^{2}$, $z$-series expansion fits are performed to parametrize the form factors over the full kinematic range. We also provide the prediction for physical observables in the rare decay $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$, including the differential branching fraction, dilepton longitudinal polarization fraction, and forward-backward asymmetries (lepton-side, hadron-side, and combined lepton-hadron). Our obtained form factors are consistent with those in other theoretical methods within the uncertainties.
@@
In the standard model (SM), flavor-changing neutral currents (FCNCs) such as the $b\to s$ transition are strongly suppressed by the Glashow-Iliopoulos-Maiani (GIM) mechanism Glashow:1970gm , as they can only proceed through loop diagrams. However, new physics scenarios could potentially enhance these FCNC processes by several orders of magnitude. Making FCNC decays excellent probes for testing the SM and searching for new physics. The LHCb Collaboration has observed several deviations between theoretical expectations and experimental measurements in $b\to sl^{+}l^{-}$ transitions, including $B^{+}\to K^{(*)+}\mu^{+}\mu^{-}$, $B^{0}\to K^{0}\mu^{+}\mu^{-}$ LHCb:2014cxe , $B^{0}\to K^{*0}\mu^{+}\mu^{-}$ LHCb:2016ykl , $B^{0}_{s}\to\phi\mu^{+}\mu^{-}$ LHCb:2021zwz , and the baryonic decay $\Lambda_{b}^{0}\to\Lambda\mu^{+}\mu^{-}$ LHCb:2015tgy . These observed anomalies motivate our theoretical investigation on the transition form factors.
While rare semileptonic decays of B mesons have been extensively studied, the corresponding baryonic decays remain less explored. This work focuses specifically on the $\Lambda_{b}\to\Lambda$ transition, aiming to provide complementary information to B meson decays. The heart of the $\Lambda_{b}\to\Lambda$ transition lies the accurate determination of the form factors, which play a pivotal role in theoretical predictions, such as the differential branching fractions of semileptonic decays and the CP-related observables in non-leptonic processes. Various theoretical approaches have been employed to investigate the $\Lambda_{b}\to\Lambda$ form factors, including soft-collinear effective theory Feldmann:2011xf , lattice QCD calculations Detmold:2012vy ; Detmold:2016pkz , and QCD light-cone sum rules Wang:2015ndk , each offering distinct advantages in addressing the challenging problem.
Among various theoretical approaches, the perturbative QCD (PQCD) approach, preserving the transverse momentum based on the $k_{T}$ factorization, has proven effective for studying heavy-to-light hadronic transitions in the fast recoil region Kurimoto:2001zj ; Li:2012nk . This approach is especially valuable as it can evaluate nonfactorizable contributions unambiguously, such as the $W$-exchange and penguin-annihilation amplitudes, which play a crucial role in predicting physical observables for nonleptonic decays, particularly the CP violation (CPV) Lu:2000em ; Cheng:2014rfa . Thus, the PQCD approach has been successfully employed to predict the CP violations of the $B^{0}\to K^{+}\pi^{-}$ and $B^{0}\to\pi^{+}\pi^{-}$ decays Belle:2004mad ; BaBar:2004gyj ; Belle:2004nch , and recently explain the long-standing puzzle of why the measured direct CP violations in $\Lambda_{b}\to p\pi^{-},pK^{-}$ are so small compared to their $B$ meson counterparts LHCb:2018fly ; LHCb:2024iis ; Han:2024kgz ; Han:2025tvc .
@@
Recent measurements have significantly advanced our understanding of CP violation in $\Lambda_{b}$ baryon decays. The LHCb collaboration firstly observed baryon CP violation in the four-body decay $\Lambda_{b}\to pK^{-}\pi^{+}\pi^{-}$, and reported a large local CP asymmetry exceeding $6.0\sigma$ significance in the region of $m_{p\pi^{+}\pi^{-}}<2.7\text{GeV}/c^{2}$ LHCb:2025ray . It implies the importance of dynamical studies of $\Lambda_{b}$ decays by the consistency between the experimental result in LHCb:2025ray and the theoretical prediction within the $N\pi$ rescattering mechanism Wang:2024oyi . Besides, the interference between resonances is instrumental in generating CP violation for multibody decays Wang:2024oyi ; Yu:2025ekh . The dominant decay modes of $\Lambda$ baryon to nucleon and pion provide substantial strong phases that may enhance local CP violation in cascade processes. The LHCb Collaboration also found evidence of CP violation in the three-body decay $\Lambda_{b}\to\Lambda K^{+}K^{-}$ LHCb:2024yzj . These experimental breakthroughs motivate our systematic investigation of the $\Lambda_{b}\to\Lambda$ transition form factors within the PQCD framework, aiming to establish a robust theoretical foundation for interpreting these intriguing observations in the future works.
@@
The first application of the PQCD approach to form factor calculations was carried out in Refs. Li:1992ce ; Kundu:1998gv , where the proton electromagnetic form factors were computed, incorporating Sudakov suppression. Subsequent studies in Refs. Shih:1998pb ; Lu:2009cm extended this approach to the hadronic transition $\Lambda_{b}\to p$, finding that the resulting form factors were significantly smaller than those obtained from nonperturbative methods. Recently, Ref. Han:2022srw confirmed the validity of the PQCD approach by including subleading-twist baryon LCDAs, successfully reproducing the $\Lambda_{b}\to p$ form factors with a reasonable magnitude.
@@
The form factors of the VAV-A current for the $\Lambda_{b}\to\Lambda$ transition were studied in Ref. Rui:2022sdc , where different parameter sets for the $\Lambda$ baryon LCDAs were examined. The parameters derived from lattice QCD Bali:2015ykx ; RQCD:2019hps only account for leading-twist LCDAs, leading to excessively small form factors. Another choice is from the light-cone QCD sum rules Liu:2014uha , which incorporates high-power corrections up to twist-6, including next-to-leading-order conformal spin expansion contributions. However, the resulting form factors are anomalously large and violate heavy-quark symmetry, suggesting that the $\Lambda$ baryon LCDA calculations given in Ref. Liu:2014uha may contain problematic assumptions.
@@
In this work, we systematically investigate all $10$ form factors describing the $\Lambda_{b}\to\Lambda$ transition induced by the $b\to s$ vector, axial-vector, and tensor currents in the PQCD approach. The $\Lambda$ baryon LCDAs with high-power corrections up to twist-6 but with leading order conformal spin expansion accuracy Liu:2009mb are employed to incorporate subleading power contributions. Our calculations successfully reproduce reasonable $\Lambda_{b}\to\Lambda$ form factors, in agreement with those from other theoretical approaches. We also demonstrate the subleading power contributions dominate baryon transition form factors, which is consistent with the conclusion in Ref. Han:2022srw .
@@
The remainder of this paper is organized as follows. Sect. II summarizes the definitions of the $\Lambda_{b}\to\Lambda$ transition form factors and presents the PQCD framework for baryon transition calculations. The LCDAs for the $\Lambda_{b}$ and $\Lambda$ baryons used in our analysis are detailed in Sect. III. We then discuss our numerical results for the form factors and their extrapolation using the modified z-expansion in Sect. IV. Sect. V presents our predictions for the differential branching fraction and angular observables in $\Lambda_{b}\to\Lambda(\to p\pi^{-})\mu^{+}\mu^{-}$ decays. Finally, we conclude with a summary in Sect. VI. The Appendix provides the explicit factorization formulas for each relevant diagram.
@@ II Definition Of The Form Factors and PQCD framework
with $p$ being the momentum of the $\Lambda_{b}$ baryon and $p^{\prime}$ the momentum of the $\Lambda$ baryon. $q=p-p^{\prime}$ denotes the momentum transferred. $\sigma^{\mu\nu}=\frac{i}{2}(\gamma^{\mu}\gamma^{\nu}-\gamma^{\nu}\gamma^{\mu})$ and $s_{\pm}=(m_{\Lambda_{b}}\pm m_{\Lambda})^{2}-q^{2}$. In some other analysis, the above matrix elements are usually decomposed into the first and second class form factors according to the Weinberg classification El-Khadra:2001wco ; Gutsche:2013pp :
@@
The relation between these two set form factors are given by Detmold:2016pkz ,
@@
In the framework of PQCD approach, the transition matrix elements of the $\Lambda_{b}\to\Lambda$ can be evaluated as the convolution of the hard kernels and baryon LCDAs,
@@
where $\Psi_{\Lambda_{b}}$ and $\Psi_{\Lambda}$ are the LCDAs of the $\Lambda_{b}$ baryon and $\Lambda$ baryon, respectively, which are non-perturbative and serve as the main sources of error in our calculation. The hard kernel $\mathcal{H}$ is derived perturbatively from diagrams. $x_{i}$ and $x^{{}^{\prime}}_{i}$ are the momentum fractions carried by quarks of $\Lambda_{b}$ and $\Lambda$, $b_{i}\left(b^{{}^{\prime}}_{i}\right)$ are the conjugate variables of the quarks’ transverse momentum $k_{iT}\left(k^{{}^{\prime}}_{iT}\right)$, respectively. The hard scale $t$ denotes the maximal virtual scale of the internal propagator within the hard kernel. Refer to Li:2001ay ; Lu:2000hj ; Han:2022srw for more details of the PQCD framework and its application in baryon decays.
@@
where $M_{\Lambda_{b}}$ and $m_{\Lambda}$ denote the masses of the $\Lambda_{b}$ and $\Lambda$ baryons, respectively. The partonic momentum with the transverse momentum retained for the quarks are parameterized as,
@@
$k_{3}$ $=\left(0,\frac{x_{3}m_{\Lambda_{b}}}{\sqrt{2}},\vec{\textbf{k}}_{3T}\right),$ $k_{3}^{\prime}$ $=\left(x_{3}^{\prime}\frac{\eta_{1}m_{\Lambda_{b}}}{\sqrt{2}},0,\vec{\textbf{k}}_{3T}^{\prime}\right),$ (22)
$k_{1}$ is the momentum of $b$ quark, $k_{1}^{{}^{\prime}}$ is the momentum of $s$ quark, and $k_{2}$ ($k_{3}$), $k_{2}^{\prime}$ ($k_{3}^{\prime}$) are the spectator $u$ ($d$) quark’s momentum in the $\Lambda_{b}$ and the $\Lambda$ baryons, respectively. $x_{i}^{(\prime)}$ and $k_{T}^{(\prime)}$ denote the corresponding momentum fractions and the the transverse momentum. In the $\Lambda_{b}$ baryon, $b$ quark carries most of the momentum fraction, $x_{1}\sim{\cal O}\left(m_{b}^{2}/M^{2}_{\Lambda_{b}}\right)$, where $m_{b}$ is the mass of b quark. $u,d$ quarks carry small momentum fractions, being the soft quarks that require at least two hard gluons to transmit momentum to them in order to form energetic quarks in the $\Lambda$ baryon, as shown in Fig. 1. In other words, ${\cal O}\left(\alpha_{s}^{2}\right)$ is the leading order for calculating the transition amplitude of $\Lambda_{b}\to\Lambda$ in the PQCD approach.
Figure 1: The leading-order Feynman diagrams for the $\Lambda_{b}\to\Lambda$ transition form factors in the PQCD approach. The green $\otimes$ denote the weak interaction currents. These diagrams are labelled by $d_{1},d_{2},\dots,d_{16}$.
@@ III Light-cone distribution amplitudes
As stated before, the necessary inputs in the PQCD approach to calculate the $\Lambda_{b}\to\Lambda$ transition form factors are the baryon LCDAs. The LCDAs of $\Lambda_{b}$ baryon are defined by the hadron matrix elements of the non-local operators sandwiched between the vacuum state and the $\Lambda_{b}$ baryon state. The general Lorentz structures in the momentum-space are Wang:2015ndk ; Bell:2013tfa ; Ball:2008fw
$\begin{split}(Y_{\Lambda_{b}})_{\alpha\beta\gamma}(x_{i},\mu)&\equiv\frac{1}{2\sqrt{2}N_{c}}\int\prod_{l=2}^{3}\frac{dz_{l}^{-}d\bm{z}_{l}}{(2\pi)^{3}}e^{ik_{l}\cdot z_{l}}\epsilon^{ijk}\langle 0|T[b_{\alpha}^{i}(0)u_{\beta}^{j}(z_{2})d_{\gamma}^{k}(z_{3})]|\Lambda_{b}\rangle\\ &=\frac{1}{8\sqrt{2}N_{c}}\Big{\{}f_{\Lambda_{b}}^{(1)}(\mu)[M_{1}(x_{2},x_{3})\gamma_{5}C^{T}]_{\gamma\beta}+f_{\Lambda_{b}}^{(2)}(\mu)[M_{2}(x_{2},x_{3})\gamma_{5}C^{T}]_{\gamma\beta}\Big{\}}[\Lambda_{b}(p)]_{\alpha},\end{split}$ (23)
where $\alpha,\beta,\gamma$ are the spinor indices, $i,j,k$ are the color indices and $N_{c}$ is the number of colors. $C^{T}$ represents the transpose of the charge conjugate matrix, and $\Lambda_{b}\left(p\right)$ represrnts the spinor of the $\Lambda_{b}$ baryon. $f_{\Lambda_{b}}^{\left(1\right)}$ and $f_{\Lambda_{b}}^{\left(2\right)}$ are normalized constants, here we choose $f_{\Lambda_{b}}^{\left(1\right)}\approx f_{\Lambda_{b}}^{\left(2\right)}\equiv f_{\Lambda_{b}}=0.03\pm 0.005GeV^{3}$ Han:2022srw ; Rui:2022sdc . The remaining parts of the chiral-even projector $M_{1}$ and the chiral-odd projector $M_{2}$ are expressed as,
@@
where $n$ and $\bar{n}$ are two light-cone vectors, $n=\left(1,0,\vec{\textbf{0}}\right)$, $\bar{n}=\left(0,1,\vec{\textbf{0}}\right)$, satisfying $n\cdot\bar{n}=1$. Several theoretical models for the $\Lambda_{b}$ baryon LCDA functions $\Psi_{2},\Psi_{3}^{+-},\Psi_{3}^{-+},\Psi_{4}$ have been proposed in Refs. Ball:2008fw ; Bell:2013tfa ; Ali:2012zza . In this work we adopt the exponential mode from Bell:2013tfa which is in a simple form and considered to be good enough in current stage Han:2022srw ; Rui:2022sdc . The exponential mode specific as follows Bell:2013tfa
@@
where the parameter $\omega_{0}$ denotes the average energy of the two light quarks and is chosen to be $0.7\pm 0.1$GeV in the PQCD approach as suggested by the calculation of the $\Lambda_{b}\to p\pi^{-},pK^{-}$ decays Han:2024kgz ; Han:2025tvc .
In the SU(3) flavor symmetry limit, the LCDAs of the final-state $\Lambda$ baryon up to twist-6 have been defined through the corresponding momentum-space projector in Liu:2009mb ; Liu:2014uha ,

### v2

@@ Abstract
We perform a systematic analysis of the $\Lambda_{b}\to\Lambda$ transition form factors using the perturbative QCD (PQCD) approach, taking into account contributions from higher-twist light-cone distribution amplitudes (LCDAs). Using inputs from lattice QCD, we show that the baryon higher-twist LCDAs give numerically dominant contributions to the form factors. By combining our PQCD results at low-$q^{2}$ region with lattice QCD predictions at high $q^{2}$, we carry out $z$-series expansion fits to obtain a unified description of the form factors over the full physical kinematic range. We also provide predictions for physical observables in the rare decay $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$, such as the differential branching fraction, the longitudinal polarization fraction of the dimuon system, and the forward–backward asymmetries.
@@
In the Standard Model (SM), processes induced by flavor-changing neutral currents (FCNCs), such as the $b\to s$ transition, are strongly suppressed by the Glashow–Iliopoulos–Maiani (GIM) mechanism Glashow:1970gm since they proceed only through loop diagrams. These FCNC decays thus serve as sensitive probes of the SM, as potential new physics contributions could alter their observable signatures. Several deviations between theoretical predictions and experimental measurements have been reported by the LHCb Collaboration in $b\to sl^{+}l^{-}$ transitions, including the decays $B^{+}\to K^{(*)+}\mu^{+}\mu^{-}$, $B^{0}\to K^{0}\mu^{+}\mu^{-}$ LHCb:2014cxe , $B^{0}\to K^{*0}\mu^{+}\mu^{-}$ LHCb:2016ykl , $B^{0}_{s}\to\phi\mu^{+}\mu^{-}$ LHCb:2021zwz , and the baryonic mode $\Lambda_{b}^{0}\to\Lambda^{0}\mu^{+}\mu^{-}$ LHCb:2015tgy . Such anomalies motivate detailed theoretical studies of the underlying transition form factors.
@@
The accurate determination of form factors is also essential in predicting CP-violating observables in nonleptonic modes. The LHCb collaboration has recently reported the first observation of CP violation in $\Lambda_{b}\to pK^{-}\pi^{+}\pi^{-}$ decay and reported a large local CP asymmetry exceeding $6.0\sigma$ significance in the region of $m_{p\pi^{+}\pi^{-}}<2.7\text{GeV}/c^{2}$ LHCb:2025ray , which significantly advance our understanding of CP violation in baryon decays. The LHCb measurements implie the importance of dynamical studies of $\Lambda_{b}$ decays by the consistency between the experimental results LHCb:2025ray and the theoretical predictions within the $N\pi$ rescattering mechanism Wang:2024oyi . Dominant decay modes of $\Lambda$ baryon to nucleon and pion provide substantial strong phases that may enhance local CP violation, as the interference between resonances is known to play an important role in generating large local CP violation Wang:2024oyi ; Yu:2025ekh . The LHCb Collaboration has also found evidence of CP violation in the three-body decay $\Lambda_{b}\to\Lambda K^{+}K^{-}$ LHCb:2024yzj . These experimental breakthroughs motivate our systematic investigation of the $\Lambda_{b}\to\Lambda$ transition form factors, aiming to establish a robust theoretical foundation for interpreting these intriguing observations in the future works.
Several theoretical approaches have been applied to calculate the $\Lambda_{b}\to\Lambda$ form factors, including the soft-collinear effective theory Feldmann:2011xf , the lattice QCD Detmold:2012vy ; Detmold:2016pkz , and the light-cone sum rules Wang:2015ndk , each offering distinct advantages in describing the form factors. Among various theoretical approaches, the perturbative QCD (PQCD) approach, preserving the transverse momentum based on the $k_{T}$ factorization, has proven effective in studying heavy-to-light hadronic transitions at the fast recoil region Kurimoto:2001zj ; Li:2012nk . The PQCD approach is especially valuable as it can calculate nonfactorizable contributions unambiguously, such as the $W$-exchange and penguin-annihilation amplitudes, which are essential for studying CP-violating observables in nonleptonic decays Lu:2000em ; Cheng:2014rfa . Thus, the PQCD approach has been successfully applied in predicting the CP violations of $B$ meson decays Belle:2004mad ; BaBar:2004gyj ; Belle:2004nch , and recently explaining the long-standing puzzle of why the measured CP violations in $\Lambda_{b}\to p\pi^{-},pK^{-}$ are so small compared to their $B$ meson counterparts LHCb:2018fly ; LHCb:2024iis ; Han:2024kgz ; Han:2025tvc .
@@
The first application of the PQCD approach to form factor calculations was carried out in Li:1992ce ; Kundu:1998gv , where the proton electromagnetic form factors were evaluated with Sudakov suppression. Subsequent works Shih:1998pb ; Lu:2009cm extended this framework to the $\Lambda_{b}\to p$ transition, though the predicted form factors were notably smaller than those from nonperturbative methods. More recently, the inclusion of high-twist baryon LCDAs in Han:2022srw has helped reconcile this discrepancy, reproducing $\Lambda_{b}\to p$ form factors of reasonable magnitude and thereby reinforcing the validity of the PQCD approach.
@@
The form factors of the VAV-A current for the $\Lambda_{b}\to\Lambda$ transition were studied in Rui:2022sdc , where different models for the $\Lambda$ baryon LCDAs were examined. The model derived from lattice QCD Bali:2015ykx ; RQCD:2019hps only account for leading-twist LCDAs, leading to excessively small form factors. Another choice for the $\Lambda$ baryon LCDAs is from QCD sum rule (QCDSR) Liu:2014uha , which includes high-power corrections up to twist-6 and next-to-leading-order conformal spin expansion contributions. However, the resulting form factors are anomalously large and violate heavy-quark symmetry Rui:2022sdc , suggesting that the $\Lambda$ baryon LCDAs obtained in Liu:2014uha by QCDSR may contain problematic assumptions.
@@
In this work, we systematically investigate all ten form factors describing the $\Lambda_{b}\to\Lambda$ transition induced by the $b\to s$ vector, axial-vector, and tensor currents within the PQCD approach. Our calculations successfully reproduce $\Lambda_{b}\to\Lambda$ form factors with reasonable magnitude, consistent with those obtained from other theoretical methods. We further show that high-twist baryon LCDAs provide the dominant contribution to baryonic transition form factors, which aligns with the conclusion drawn in Han:2022srw .
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the definitions of the $\Lambda_{b}\to\Lambda$ transition form factors and presents the PQCD framework for baryonic transition form factors. The LCDAs for the $\Lambda_{b}$ and $\Lambda$ baryons used in this work are listed in Sec. III in detail. Numerical results for the form factors are presented in Sec. IV, where we also perform extrapolations using the modified $z$-expansion. Our predictions for the differential branching fraction and angular observables in $\Lambda_{b}\to\Lambda(\to p\pi^{-})\mu^{+}\mu^{-}$ decays are given in Sec. V. Finally, we conclude with a summary in Sec. VI. The Appendix provides the explicit factorization formulas for each relevant diagram.
@@ II Definition Of The Form Factors and PQCD framework
with $p$ being the momentum of the $\Lambda_{b}$ baryon and $p^{\prime}$ the momentum of the $\Lambda$ baryon. $q=p-p^{\prime}$ denotes the momentum transferred. $\sigma^{\mu\nu}\equiv\frac{i}{2}(\gamma^{\mu}\gamma^{\nu}-\gamma^{\nu}\gamma^{\mu})$ and $s_{\pm}\equiv(m_{\Lambda_{b}}\pm m_{\Lambda})^{2}-q^{2}$. The above matrix elements are also usually decomposed into the first and second class form factors according to the Weinberg classification El-Khadra:2001wco ; Gutsche:2013pp ,
@@
The relations between these two definitions are derived as Detmold:2016pkz ,
@@
In the framework of PQCD approach, the $\Lambda_{b}\to\Lambda$ transition matrix elements can be expressed as convolution of the hard kernel with the baryon LCDAs,
@@
where $\Psi_{\Lambda_{b}}$ and $\Psi_{\Lambda}$ denote the LCDAs of the $\Lambda_{b}$ and $\Lambda$ baryons, respectively. These wave functions are nonperturbative and constitute the dominant source of theoretical uncertainty in our calculation. The hard kernel $\mathcal{H}$ is computed perturbatively from the relevant Feynman diagrams. $x_{i}$ and $x^{{}^{\prime}}_{i}$ represent the longitudinal momentum fractions carried by quarks inside the $\Lambda_{b}$ and $\Lambda$, while $b_{i}\left(b^{{}^{\prime}}_{i}\right)$ are the conjugate variables to the quarks’ transverse momentum $k_{iT}\left(k^{{}^{\prime}}_{iT}\right)$, respectively. The hard scale $t$ corresponds to the maximum virtuality of internal propagators in $\mathcal{H}$. Further details on the PQCD formalism and its application to baryon decays can be found in Li:2001ay ; Lu:2000hj ; Han:2022srw .
@@
where $M_{\Lambda_{b}}$ and $m_{\Lambda}$ denote the masses of the $\Lambda_{b}$ and $\Lambda$, respectively. The momentum for quarks with the transverse momentum retained are parameterized as,
@@
$k_{3}$ $=\left(0,\frac{x_{3}m_{\Lambda_{b}}}{\sqrt{2}},\vec{\textbf{k}}_{3T}\right),$ $k_{3}^{\prime}$ $=\left(x_{3}^{\prime}\frac{\eta_{1}m_{\Lambda_{b}}}{\sqrt{2}},0,\vec{\textbf{k}}_{3T}^{\prime}\right).$ (22)
$k_{1}$ and $k_{1}^{\prime}$ represent the momenta of the $b$ and $s$ quarks, respectively, while $k_{2}$ ($k_{3}$) and $k_{2}^{\prime}$ ($k_{3}^{\prime}$) denote the momenta of the spectator $u$ ($d$) quarks, respectively. In the $\Lambda_{b}$ baryon, $b$ quark carries the majority of the longitudinal momentum, $x_{1}\sim{\cal O}\left(m_{b}^{2}/M^{2}_{\Lambda_{b}}\right)$, where $m_{b}$ is the mass of b quark. $u,d$ quarks carry small momentum and are treated as soft degrees of freedom. As illustrated in Fig. 1, at least two hard gluon exchanges are required to transfer momentum to these soft quarks to form the energetic $\Lambda$ baryon in the final state. This implies that the $\Lambda_{b}\to\Lambda$ transition amplitude in the PQCD approach starts at $\mathcal{O}(\alpha_{s}^{2})$, which constitutes the leading-order contribution.
Figure 1: The leading order Feynman diagrams for the $\Lambda_{b}\to\Lambda$ transition form factors in the PQCD approach. The green $\otimes$ denote the weak interaction currents.
@@ III Light-cone distribution amplitudes
As stated earlier, the essential inputs for calculating the $\Lambda_{b}\to\Lambda$ transition form factors in the PQCD approach are the LCDAs of the baryons. The LCDAs of the $\Lambda_{b}$ baryon are defined by the matrix elements of non-local operators between the vacuum and the $\Lambda_{b}$ baryon state. The corresponding Lorentz structures in momentum space take the form Wang:2015ndk ; Bell:2013tfa ; Ball:2008fw
$\begin{split}(Y_{\Lambda_{b}})_{\alpha\beta\gamma}(x_{i},\mu)&\equiv\frac{1}{2N_{c}}\int\prod_{l=2}^{3}\frac{dz_{l}^{-}d\bm{z}_{l}}{(2\pi)^{3}}e^{ik_{l}\cdot z_{l}}\epsilon^{ijk}\langle 0|T[b_{\alpha}^{i}(0)u_{\beta}^{j}(z_{2})d_{\gamma}^{k}(z_{3})]|\Lambda_{b}\rangle\\ &=\frac{1}{8N_{c}}\Big\{f_{\Lambda_{b}}^{(1)}(\mu)[M_{1}(x_{2},x_{3})\gamma_{5}C^{T}]_{\gamma\beta}+f_{\Lambda_{b}}^{(2)}(\mu)[M_{2}(x_{2},x_{3})\gamma_{5}C^{T}]_{\gamma\beta}\Big\}[\Lambda_{b}(p)]_{\alpha},\end{split}$ (23)
where $\alpha,\beta,\gamma$ are the spinor indices, $i,j,k$ are the color indices and $N_{c}$ is the number of colors. $C^{T}$ represents the transpose of the charge conjugate matrix, and $\Lambda_{b}\left(p\right)$ represrnts the spinor of the $\Lambda_{b}$ baryon. $f_{\Lambda_{b}}^{\left(1\right)}$ and $f_{\Lambda_{b}}^{\left(2\right)}$ are normalized constants, here we choose $f_{\Lambda_{b}}^{\left(1\right)}\approx f_{\Lambda_{b}}^{\left(2\right)}\equiv f_{\Lambda_{b}}=0.031\pm 0.005GeV^{3}$ Han:2022srw ; Rui:2022sdc . The remaining parts of the chiral-even projector $M_{1}$ and the chiral-odd projector $M_{2}$ are expressed as,
@@
where $n$ and $\bar{n}$ are two light-cone vectors, $n=\left(1,0,\vec{\textbf{0}}_{T}\right)$, $\bar{n}=\left(0,1,\vec{\textbf{0}}_{T}\right)$, satisfying $n\cdot\bar{n}=1$. Several theoretical models for the $\Lambda_{b}$ baryon LCDA functions $\Psi_{2},\Psi_{3}^{+-},\Psi_{3}^{-+},\Psi_{4}$ have been proposed in Ball:2008fw ; Bell:2013tfa ; Ali:2012zza . In this work we adopt the exponential mode from Bell:2013tfa which is in a simple form and considered to be good enough in current stage Han:2022srw ; Rui:2022sdc . The exponential mode is given by Bell:2013tfa
@@
where the parameter $\omega_{0}$ denotes the average energy of the two light quarks and is chosen to be $0.7\pm 0.1$GeV in the PQCD approach as suggested by the analyses of the $\Lambda_{b}\to p\pi^{-},pK^{-}$ decays Han:2024kgz ; Han:2025tvc .
In the SU(3) flavor symmetry limit, the LCDAs for an outgoing final-state $\Lambda$ baryon up to twist-6 have been defined through the corresponding momentum-space projector in Braun:2000kw ; Liu:2014uha ; Wein:2015oqa ,
$(\overline{Y}_{\Lambda})_{\alpha\beta\gamma}(x_{i}^{\prime},\mu)=$ $\frac{1}{2\sqrt{2}N_{c}}\int\prod_{l=2}^{3}\frac{dz_{l}^{-}d\bm{z}_{l}}{(2\pi)^{3}}e^{ik_{l}\cdot z_{l}}\epsilon^{ijk}\langle\Lambda(p^{\prime})|\epsilon^{ijk}{\bar{u}}_{\alpha}^{i}(0){\bar{d}}_{\beta}^{j}(z_{2}){\bar{s}}_{\gamma}^{k}(z_{3})|0\rangle\$
$=$ $\frac{-1}{8\sqrt{2}N_{c}}\Big\{S_{1}m_{\Lambda}C_{\beta\alpha}(\bar{\Lambda}^{+}\gamma_{5})_{\gamma}+S_{2}m_{\Lambda}C_{\beta\alpha}(\bar{\Lambda}^{-}\gamma_{5})_{\gamma}+P_{1}m_{\Lambda}(C\gamma_{5})_{\beta\alpha}\bar{\Lambda}^{+}_{\gamma}$
@@
$+V_{3}\frac{m_{\Lambda}}{2}(C\gamma_{\perp})_{\beta\alpha}(\bar{\Lambda}^{+}\gamma_{5}\gamma^{\perp})_{\gamma}+V_{4}\frac{m_{\Lambda}}{2}(C\gamma_{\perp})_{\beta\alpha}(\bar{\Lambda}^{-}\gamma_{5}\gamma^{\perp})_{\gamma}$
[…]
$-T_{2}(iC\sigma_{\perp\mathcal{P}^{\prime}})_{\beta\alpha}(\bar{\Lambda}^{-}\gamma_{5}\gamma^{\perp})_{\gamma}-T_{3}\frac{m_{\Lambda}}{\mathcal{P}^{\prime}z}(iC\sigma_{\mathcal{P}^{\prime}z})_{\beta\alpha}(\bar{\Lambda}^{+}\gamma_{5})_{\gamma}$
$-T_{4}\frac{m_{\Lambda}}{\mathcal{P}^{\prime}z}(iC\sigma_{z\mathcal{P}^{\prime}})_{\beta\alpha}(\bar{\Lambda}^{-}\gamma_{5})_{\gamma}-T_{5}\frac{m_{\Lambda}^{2}}{2\mathcal{P}^{\prime}z}(iC\sigma_{\perp z})_{\beta\alpha}(\bar{\Lambda}^{+}\gamma_{5}\gamma^{\perp})_{\gamma}$
$-T_{6}\frac{m_{\Lambda}^{2}}{2\mathcal{P}^{\prime}z}(iC\sigma_{\perp z})_{\beta\alpha}(\bar{\Lambda}^{-}\gamma_{5}\gamma^{\perp})_{\gamma}+T_{7}\frac{m_{\Lambda}}{2}(C\sigma_{\perp\perp^{\prime}})_{\beta\alpha}(\bar{\Lambda}^{+}\gamma_{5}\sigma^{\perp\perp^{\prime}})_{\gamma}$
$+T_{8}\frac{m_{\Lambda}}{2}(C\sigma_{\perp\perp^{\prime}})_{\beta\alpha}(\bar{\Lambda}^{-}\gamma_{5}\sigma^{\perp\perp^{\prime}})_{\gamma}\Big\},$ (26)